# Dense Hopfield Networks in the Teacher-Student Setting

### v1

@@ Appendix A Gardner’s Hamiltonian vs Krotov’s Hamiltonian
On the other hand, Krotov’s Hamiltonian may be rewritten
@@
where the sum over $i_{1}...i_{p}$ includes both the set of indices $i_{1}\neq...\neq i_{p}$ found in $H\left[\sigma\right]$ and other configurations where some indices are equal. For example, the configuration $i_{1}\neq...\neq i_{p-1}=i_{p}$ contains the fewest equal indices after $i_{1}\neq...\neq i_{p}$. In other words, $\mathcal{H}\left[\sigma\right]$ can be expressed as an expansion around $H\left[\sigma\right]$, and the two Hamiltonians are equivalent when the normalized residuals $\frac{\mathcal{H}\left[\sigma\right]-H\left[\sigma\right]}{N}$ vanish in the limit of large $N$. In this study, we will encounter two cases which will bring different results.
@@
We use the expansion of $\mathcal{H}\left[\sigma\right]$ to discuss both the aligned case and the misaligned case. We start by writing the $i_{1}\neq...\neq i_{p}$ and $i_{1}\neq...\neq i_{p-1}=i_{p}$ terms explicitly, which leads to the form:
@@
$\quad-\frac{1}{2}\frac{p(p-1)}{N^{p-1}}\sum_{\mu}\sum_{i_{1}\neq...\neq i_{p-1}}\xi_{i_{1}}^{\mu}...\left(\xi_{i_{p-1}}^{\mu}\right)^{2}\sigma_{i_{1}}...\left(\sigma_{i_{p-1}}\right)^{2}+...$
@@
In the aligned case, $H^{\prime}\left[\sigma\right]$ is $\mathcal{O}\left(1\right)$ in $N$ because the sum over $i_{1}<...<i_{p-2}$ is $\mathcal{O}\left(N^{p-2}\right)$. The terms implied by the ellipsis are even smaller because their sums are resctricted by more equality constraints. Therefore, the residuals $\frac{\mathcal{H}\left[\sigma\right]-H\left[\sigma\right]}{N}$ vanish in the limit of large $N$, and the two Hamiltonians are equivalent. Conversely, we find that $\mathcal{H}\left[\sigma\right]$ and $H\left[\sigma\right]$ differ from each other in the misaligned case (see Appendix B for more details). Therefore, although the phases of $H\left[\sigma\right]$ that we obtain in this study are qualitatively similar to the ones observed by Krotov [26, 37], the phase diagram of $H\left[\sigma\right]$ must be compared against a simulation of $H\left[\sigma\right]$ rather than $\mathcal{H}\left[\sigma\right]$ in order to test our theory quantitatively.
To understand how to sample $\sigma$ in both models, consider a Monte-Carlo simulation used to find the statistical equilibrium of a spin ensemble $\sigma$ with Hamiltonian $G\left[\sigma\right]$. To be more specific, suppose $\sigma$ is updated to a randomly selected new state $\sigma^{\prime}$ with acceptance probability $P_{i}=\frac{1}{1+\exp\left[\beta\left(G\left[\sigma^{\prime}\right]-G\left[\sigma\right]\right)\right]}$ for a large number of time-steps. This approach works well for $G\left[\sigma\right]=\mathcal{H}\left[\sigma\right]$. However, in the case of $H\left[\sigma\right]$, we find that the simulation only converges when we use the local field $h_{i}=\frac{p!}{N^{p-1}}\sum_{\mu}\xi^{\mu}_{i}\sum_{i_{1}<...<i_{p-1}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p-1}}\sigma_{i_{1}}...\sigma_{i_{p-1}}$ mentioned by Gardner [30] to approximate $\frac{H\left[\sigma^{\prime}\right]-H\left[\sigma\right]}{2\sigma_{i}}$ at large $N$. In other words, we iteratively flip randomly chosen spins $\sigma_{i}$ with acceptance probability $P_{i}=\frac{1}{1+\exp\left(2\beta h_{i}\sigma_{i}\right)}$ for a large number of time steps. For arbitrary $p$, it is not obvious how to compute $h_{i}$ quickly as a sub-routine of the Monte-Carlo simulation. However, we find that both $p=3$ and $p=4$ have closed-formed expressions that are easy to evaluate numerically in an efficient way. To be more precise,
@@
The output of the neural network model that Krotov designed for classification of data is $c_{j}=\tanh\left[\frac{1}{2}\beta\left(\mathcal{H}\left[\sigma^{\prime}\right]-\mathcal{H}\left[\sigma\right]\right)\right]\approx\tanh\left[\beta p\sum_{\mu}\xi^{\mu}_{j}\left(\frac{1}{N}\sum_{i}\xi^{\mu}_{i}\sigma_{i}\right)^{p-1}\right]$. We omit the linear rectifier present in the original paper [26] because the overlaps $\frac{1}{N}\sum_{i}\xi_{i}^{\mu}\sigma_{i}$ are almost always positive (see for example the Supplement of [72]). The predicted class is then $j^{\prime}=\operatorname*{argmax}_{j}\left\{c_{i}\right\}$. Using $1-P_{j}=\frac{1}{1+\exp\left[\beta\left(\mathcal{H}\left[\sigma^{\prime}\right]-\mathcal{H}\left[\sigma\right]\right)\right]}$ instead of $c_{j}$ does not change $j^{\prime}$ because $1-P_{j}$ and $c_{j}$ are related by $1-P_{j}=\frac{1}{2}\left[c_{j}+1\right]$. When we evaluate $P_{i}$ using $H$ instead of $\mathcal{H}$, this relation does not always hold exactly. Rather, it should be considered an approximation.
@@ Appendix B Direct model cumulant expansions
$\left\langle Z^{L}\right\rangle$ $=\left\langle\sum_{\sigma}\exp\left(-\beta\sum_{\gamma=1}^{L}H\left[\sigma^{\gamma}|\xi\right]\right)\right\rangle$
with $\sigma=\begin{Bmatrix}\sigma^{1}\ \dots\ \sigma^{L}\end{Bmatrix}$. Gardner simplifies it to:
$\left\langle Z^{L}\right\rangle$ $\approx\Bigg{\langle}\sum_{\sigma}\exp\Bigg{(}\beta N\sum_{\gamma}\sum_{\mu\in\Gamma_{\gamma}}\left[\frac{1}{N}\sum_{i}\xi^{\mu}_{i}\sigma^{\gamma}_{i}\right]^{p}$ (9)
$\quad+\beta\sum_{\gamma}\sum_{\mu\in\bar{\Gamma}}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\Bigg{)}\Bigg{\rangle},$
@@
$\approx\beta\left\langle\sum_{\gamma}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\right\rangle$
$\quad+\frac{1}{2}\beta^{2}\left\langle\left[\sum_{\gamma}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\right]^{2}\right\rangle$
$\approx\frac{1}{2}\beta^{2}\left\langle\left[\sum_{\gamma}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\right]\left[\sum_{\delta}\frac{p!}{N^{p-1}}\sum_{j_{1}<...<j_{p}}\xi^{\mu}_{j_{1}}...\xi^{\mu}_{j_{p}}\sigma^{\delta}_{j_{1}}...\sigma^{\delta}_{j_{p}}\right]\right\rangle$
@@
$=\frac{1}{2}\beta^{2}\left[\frac{p!}{N^{p-1}}\right]^{2}\left\langle\sum_{\gamma}\sum_{\delta}\sum_{i_{1}<...<i_{p}}\sum_{j_{1}<...<j_{p}}\xi^{\mu}_{i_{1}}\xi^{\mu}_{j_{1}}...\xi^{\mu}_{i_{p}}\xi^{\mu}_{j_{p}}\ \sigma^{\gamma}_{i_{1}}\sigma^{\delta}_{j_{1}}...\sigma^{\gamma}_{i_{p}}\sigma^{\delta}_{j_{p}}\right\rangle.$
@@
$=\frac{1}{2}\beta^{2}\left[\frac{p!}{N^{p-1}}\right]^{2}\sum_{\gamma}\sum_{\delta}\sum_{i_{1}<...<i_{p}}\sum_{j_{1}<...<j_{p}}\delta_{i_{1}j_{1}}...\delta_{i_{p}j_{p}}\ \sigma^{\gamma}_{i_{1}}\sigma^{\delta}_{j_{1}}...\sigma^{\gamma}_{i_{p}}\sigma^{\delta}_{j_{p}}$
$=\frac{1}{2}\beta^{2}\left[\frac{p!}{N^{p-1}}\right]^{2}\sum_{\gamma}\sum_{\delta}\sum_{i_{1}<...<i_{p}}\sigma^{\gamma}_{i_{1}}\sigma^{\delta}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\sigma^{\delta}_{i_{p}}$
$\approx\frac{1}{2}\beta^{2}\frac{p!}{N^{p-1}}\frac{1}{N^{p-1}}\sum_{\gamma\delta}\left[\sum_{i}\sigma^{\gamma}_{i}\sigma^{\delta}_{i}\right]^{p}$
$=\beta^{2}\frac{p!}{N^{p-1}}N\sum_{\gamma<\delta}\left[\frac{1}{N}\sum_{i}\sigma^{\gamma}_{i}\sigma^{\delta}_{i}\right]^{p}+\frac{1}{2}\beta^{2}\frac{p!}{N^{p-1}}LN.$
The order $n>2$ terms are subdominant in $N$ and can be neglected when $p\geq 3$ [30]. The RS free entropy is then obtained through a standard approach to the replica method. Note that Gardner’s Hamiltonian is misaligned with $\sigma$ when the free entropy is dominated by this cumulant expansion (see Case 2 of Appendix A). In the case of Krotov’s Hamiltonian, we must also take into account the correction $H^{\prime}\left[\sigma\right]=\frac{1}{2}\frac{p!}{N^{p-2}}\sum_{\gamma}\sum_{i_{1}<...<i_{p-2}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p-2}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p-2}}$ introduced in appendix A by imposing $i_{p-1}=i_{p}$. In fact, a cumulant expansion of this expression gives
@@
$\approx\frac{1}{4}\beta^{2}\frac{p!}{N^{p-2}}\frac{p(p-1)}{N^{p-2}}\sum_{\gamma<\delta}\left[\sum_{i}\sigma^{\gamma}_{i}\sigma^{\delta}_{i}\right]^{p-2}+\frac{1}{8}\beta^{2}\frac{p!}{N^{p-2}}L$
$=\frac{1}{4}p(p-1)\beta^{2}\frac{p!}{N^{p-1}}N\sum_{\gamma<\delta}\left[\frac{1}{N}\sum_{i}\sigma^{\gamma}_{i}\sigma^{\delta}_{i}\right]^{p-2}+\frac{1}{8}\beta^{2}\frac{p!}{N^{p-1}}LN,$
which contributes to the free energy on the same order in $N$ as Gardner’s Hamiltonian. Therefore, Krotov’s Hamiltonian is not equivalent to Gardner’s Hamiltonian when the latter is misaligned with $\sigma$ (see Case 2). The index configurations with more equality constraints also contribute to the free entropy on the same order in $N$ because the factors of $N$ that are lost to equality constraints are restored when the sums get squared in the cumulant expansion.
$p=2$ is the only positive integer such that Gardner’s Hamiltonian and Krotov’s Hamiltonian are equivalent [5, 30]. In the misaligned case with a single stored pattern $\xi^{*}$ (see Case 2), the free entropy of $p=2$ simplifies to:

### v2

@@ Appendix A Gardner’s Hamiltonian vs K & H’s Hamiltonian
On the other hand, K & H’s Hamiltonian may be rewritten
@@
where the sum over $i_{1}...i_{p}$ includes both the set of indices $i_{1}\neq...\neq i_{p}$ found in $H\left[\sigma\right]$ and other configurations where some indices are equal. For example, the configuration $i_{1}\neq...\neq i_{p-1}=i_{p}$ contains the fewest equal indices after $i_{1}\neq...\neq i_{p}$. In other words, $\mathcal{H}\left[\sigma\right]$ can be expressed as an expansion around $H\left[\sigma\right]$, and the two Hamiltonians are equivalent when the normalized residuals $\frac{\mathcal{H}\left[\sigma\right]-H\left[\sigma\right]}{N}$ vanish in the limit of large $N$. In this study, we encounter two cases which bring different results.
@@
We use the expansion of $\mathcal{H}\left[\sigma\right]$ to discuss both the aligned case and the misaligned case. We start by writing the $i_{1}\neq...\neq i_{p}$ and $i_{1}\neq...\neq i_{p-1}=i_{p}$ terms explicitly, which leads to the form
@@
$\quad-\frac{1}{2}\frac{p(p-1)}{N^{p-1}}\sum_{\mu}\sum_{i_{1}\neq...\neq i_{p-1}}\xi_{i_{1}}^{\mu}...\left(\xi_{i_{p-1}}^{\mu}\right)^{2}\sigma_{i_{1}}...\left(\sigma_{i_{p-1}}\right)^{2}+...\,,$
@@
In the aligned case, $H^{\prime}\left[\sigma\right]$ is $\mathcal{O}\left(1\right)$ in $N$ because the sum over $i_{1}<...<i_{p-2}$ is $\mathcal{O}\left(N^{p-2}\right)$. The terms implied by the ellipsis are even smaller because their sums are resctricted by more equality constraints. Therefore, the residuals $\frac{\mathcal{H}\left[\sigma\right]-H\left[\sigma\right]}{N}$ vanish in the limit of large $N$, and the two Hamiltonians are equivalent. Conversely, we find that $\mathcal{H}\left[\sigma\right]$ and $H\left[\sigma\right]$ differ from each other in the misaligned case (see Appendix B for more details). Therefore, although the phases of $H\left[\sigma\right]$ that we obtain in this study are qualitatively similar to the ones observed by K & H [26, 37], the phase diagram of $H\left[\sigma\right]$ must be compared against a simulation of $H\left[\sigma\right]$ rather than $\mathcal{H}\left[\sigma\right]$ in order to test our theory quantitatively.
To understand how to sample $\sigma$ in both models, consider a Monte-Carlo simulation used to find the statistical equilibrium of a spin ensemble $\sigma$ with Hamiltonian $G\left[\sigma\right]$. To be more specific, suppose $\sigma$ is updated to a new state $\sigma^{\prime}$ with a randomly selected spin $\sigma_{i}$ flipped with acceptance probability $P_{i}=\frac{1}{1+\exp\left[\beta\left(G\left[\sigma^{\prime}\right]-G\left[\sigma\right]\right)\right]}$ for a large number of time-steps. This approach works well for $G\left[\sigma\right]=\mathcal{H}\left[\sigma\right]$. However, in the case of $H\left[\sigma\right]$, we find that the simulation only converges when we use the local field $h_{i}=\frac{p!}{N^{p-1}}\sum_{\mu}\xi^{\mu}_{i}\sum_{i_{1}<...<i_{p-1}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p-1}}\sigma_{i_{1}}...\sigma_{i_{p-1}}$ mentioned by Gardner [30] to approximate $\frac{H\left[\sigma^{\prime}\right]-H\left[\sigma\right]}{2\sigma_{i}}$ at large $N$. In other words, we iteratively flip randomly chosen spins $\sigma_{i}$ with acceptance probability $P_{i}=\frac{1}{1+\exp\left(2\beta h_{i}\sigma_{i}\right)}$ for a large number of time steps. For arbitrary $p$, it is not obvious how to compute $h_{i}$ quickly as a sub-routine of the Monte-Carlo simulation. However, we find that both $p=3$ and $p=4$ have closed-formed expressions that are easy to evaluate numerically in an efficient way. To be more precise,
@@
The output of the neural network model that K & H designed for classification of data is $c_{j}=\tanh\left[\frac{1}{2}\beta\left(\mathcal{H}\left[\sigma^{\prime}\right]-\mathcal{H}\left[\sigma\right]\right)\right]\approx\tanh\left[\beta p\sum_{\mu}\xi^{\mu}_{j}\left(\frac{1}{N}\sum_{i}\xi^{\mu}_{i}\sigma_{i}\right)^{p-1}\right]$. We omit the linear rectifier present in the original paper [26] because the overlaps $\frac{1}{N}\sum_{i}\xi_{i}^{\mu}\sigma_{i}$ are almost always positive (see for example the Supplement of [78]). The predicted class is then $j^{\prime}=\operatorname*{argmax}_{j}\left\{c_{i}\right\}$. Using $1-P_{j}=\frac{1}{1+\exp\left[\beta\left(\mathcal{H}\left[\sigma^{\prime}\right]-\mathcal{H}\left[\sigma\right]\right)\right]}$ instead of $c_{j}$ does not change $j^{\prime}$ because $1-P_{j}$ and $c_{j}$ are related by $1-P_{j}=\frac{1}{2}\left[c_{j}+1\right]$. When we evaluate $P_{i}$ using $H$ instead of $\mathcal{H}$, this relation does not always hold exactly. Rather, it should be considered an approximation.
@@ Appendix B Direct model cumulant expansions
$\left\langle Z^{L}\right\rangle$ $=\left\langle\sum_{\sigma}\exp\left(-\beta\sum_{\gamma=1}^{L}H\left[\sigma^{\gamma}|\xi\right]\right)\right\rangle\,,$
with $\sigma=\begin{Bmatrix}\sigma^{1}\ \dots\ \sigma^{L}\end{Bmatrix}$. Gardner simplifies it to
$\left\langle Z^{L}\right\rangle\approx\Bigg{\langle}\sum_{\sigma}\exp\Bigg{(}\beta N\sum_{\gamma}\sum_{\mu\in\Gamma_{\gamma}}\left[\frac{1}{N}\sum_{i}\xi^{\mu}_{i}\sigma^{\gamma}_{i}\right]^{p}+\beta\sum_{\gamma}\sum_{\mu\in\bar{\Gamma}}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\Bigg{)}\Bigg{\rangle}\,,$ (10)
@@
$\quad\approx\beta\left\langle\sum_{\gamma}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\right\rangle+\frac{1}{2}\beta^{2}\left\langle\left[\sum_{\gamma}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\right]^{2}\right\rangle$
$\quad\approx\frac{1}{2}\beta^{2}\left\langle\left[\sum_{\gamma}\frac{p!}{N^{p-1}}\sum_{i_{1}<...<i_{p}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\right]\left[\sum_{\delta}\frac{p!}{N^{p-1}}\sum_{j_{1}<...<j_{p}}\xi^{\mu}_{j_{1}}...\xi^{\mu}_{j_{p}}\sigma^{\delta}_{j_{1}}...\sigma^{\delta}_{j_{p}}\right]\right\rangle\,,$
@@
$\quad=\frac{1}{2}\beta^{2}\left[\frac{p!}{N^{p-1}}\right]^{2}\left\langle\sum_{\gamma}\sum_{\delta}\sum_{i_{1}<...<i_{p}}\sum_{j_{1}<...<j_{p}}\xi^{\mu}_{i_{1}}\xi^{\mu}_{j_{1}}...\xi^{\mu}_{i_{p}}\xi^{\mu}_{j_{p}}\ \sigma^{\gamma}_{i_{1}}\sigma^{\delta}_{j_{1}}...\sigma^{\gamma}_{i_{p}}\sigma^{\delta}_{j_{p}}\right\rangle\,.$
@@
$\quad=\frac{1}{2}\beta^{2}\left[\frac{p!}{N^{p-1}}\right]^{2}\sum_{\gamma}\sum_{\delta}\sum_{i_{1}<...<i_{p}}\sum_{j_{1}<...<j_{p}}\delta_{i_{1}j_{1}}...\delta_{i_{p}j_{p}}\ \sigma^{\gamma}_{i_{1}}\sigma^{\delta}_{j_{1}}...\sigma^{\gamma}_{i_{p}}\sigma^{\delta}_{j_{p}}$
$\quad=\frac{1}{2}\beta^{2}\left[\frac{p!}{N^{p-1}}\right]^{2}\sum_{\gamma}\sum_{\delta}\sum_{i_{1}<...<i_{p}}\sigma^{\gamma}_{i_{1}}\sigma^{\delta}_{i_{1}}...\sigma^{\gamma}_{i_{p}}\sigma^{\delta}_{i_{p}}$
$\quad\approx\frac{1}{2}\beta^{2}\frac{p!}{N^{p-1}}\frac{1}{N^{p-1}}\sum_{\gamma\delta}\left[\sum_{i}\sigma^{\gamma}_{i}\sigma^{\delta}_{i}\right]^{p}$
$\quad=\beta^{2}\frac{p!}{N^{p-1}}N\sum_{\gamma<\delta}\left[\frac{1}{N}\sum_{i}\sigma^{\gamma}_{i}\sigma^{\delta}_{i}\right]^{p}+\frac{1}{2}\beta^{2}\frac{p!}{N^{p-1}}LN\,.$
The order $n>2$ terms are subdominant in $N$ and can be neglected when $p\geq 3$ [30]. The RS free entropy is then obtained through a standard approach to the replica method. Note that Gardner’s Hamiltonian is misaligned with $\sigma$ when the free entropy is dominated by this cumulant expansion (see Case 2 of Appendix A). In the case of K & H’s Hamiltonian, we must also take into account the correction $H^{\prime}\left[\sigma\right]=\frac{1}{2}\frac{p!}{N^{p-2}}\sum_{\gamma}\sum_{i_{1}<...<i_{p-2}}\xi^{\mu}_{i_{1}}...\xi^{\mu}_{i_{p-2}}\sigma^{\gamma}_{i_{1}}...\sigma^{\gamma}_{i_{p-2}}$ introduced in appendix A by imposing $i_{p-1}=i_{p}$. In fact, a cumulant expansion of this expression gives
@@
$\quad\approx\frac{1}{4}\beta^{2}\frac{p!}{N^{p-2}}\frac{p(p-1)}{N^{p-2}}\sum_{\gamma<\delta}\left[\sum_{i}\sigma^{\gamma}_{i}\sigma^{\delta}_{i}\right]^{p-2}+\frac{1}{8}\beta^{2}\frac{p!}{N^{p-2}}L$
$\quad=\frac{1}{4}p(p-1)\beta^{2}\frac{p!}{N^{p-1}}N\sum_{\gamma<\delta}\left[\frac{1}{N}\sum_{i}\sigma^{\gamma}_{i}\sigma^{\delta}_{i}\right]^{p-2}+\frac{1}{8}\beta^{2}\frac{p!}{N^{p-1}}LN\,,$
which contributes to the free energy on the same order in $N$ as Gardner’s Hamiltonian. Therefore, K & H’s Hamiltonian is not equivalent to Gardner’s Hamiltonian when the latter is misaligned with $\sigma$ (see Case 2). The index configurations with more equality constraints also contribute to the free entropy on the same order in $N$ because the factors of $N$ that are lost to equality constraints are restored when the sums get squared in the cumulant expansion.
$p=2$ is the only positive integer such that Gardner’s Hamiltonian and Krotov’s Hamiltonian are equivalent [5, 30]. In the misaligned case with a single stored pattern $\xi^{*}$ (see Case 2), the free entropy of $p=2$ simplifies to